# Entangling Quantum Memories at Channel Capacity

### v1

@@ Methods
Given a state variance of $\sigma^{2}$, the probability of making $k$-shift errors in the logical measurement for sq-GKP states (along either the Pauli $X$ or $Z$ basis) is given by,
$P^{\mathrm{sq}}(X^{k},\sigma^{2})\equiv P^{\mathrm{sq}}(Z^{k},\sigma^{2})=\sum_{j\in\mathbb{Z}}\frac{1}{2}\left(\mathrm{erf}\left(\sqrt{\frac{2\pi}{d}}\frac{jd+k+1/2}{\sigma}\right)-\mathrm{erf}\left(\sqrt{\frac{2\pi}{d}}\frac{jd+k-1/2}{\sigma}\right)\right).$ (38)
where $k\in\{0,1,2,\ldots,d-1\}$, with $k=0$ denoting the probability of deducing a correct logical outcome, whereas the summation over the index $j$ accounts for shifts of magnitude greater than one lattice spacing. Assuming that Bell state measurement logical outcomes are $l_{Z},l_{X}$ the final state can be expressed as
$\tilde{\rho}_{\bm{M}_{1}\bm{M}_{2}}=\sum_{k_{1},k_{2}=0}^{d-1}P_{\mathrm{shift}}^{(k_{1},k_{2})}(\sigma^{2})\,\ket{\Psi_{k^{\prime},l^{\prime}}}\!\!\bra{\Psi_{k^{\prime},l^{\prime}}}_{\bm{M}_{1}\bm{M}_{2}}$ (39)
where $P_{\mathrm{shift}}^{(k_{1},k_{2};)}(\sigma^{2})=P(X^{k_{1}},\sigma^{2})\cdot P(Z^{k_{2}},\sigma^{2})$ and the state |$\ket{\Psi_{k^{\prime},l^{\prime}}}\equiv W_{1}^{(0,d-l_{Z})}W_{2}^{(d-l_{X},0)}\ket{\Psi_{0,0}}$ is a different qudit Bell state. The hashing bound [22] based lower bound to the distillable entanglement for this state is given by
@@
$\tilde{P}_{\mathrm{shift}}(X^{k};1-\eta)\approx\frac{1}{2}\left(\frac{\left|k+1/2\right|}{k+1/2}-\frac{\left|k-1/2\right|}{k-1/2}+\frac{\sqrt{d\varepsilon}\exp\left(-2\pi(k-1/2)^{2}/(d\varepsilon)\right)}{\sqrt{2}\pi(k-1/2)}-\frac{\sqrt{d\varepsilon}\exp\left(-2\pi(k+1/2)^{2}/(d\varepsilon)\right)}{\sqrt{2}\pi(k+1/2)}\right)$ (41)
@@
$\mathrm{erf}\left(\frac{x-\mu}{\sigma}\right)=\frac{\left(\frac{\sigma}{\sqrt{\pi}}+\mathcal{\mathcal{O}}\left(\sigma^{2}\right)\right)e^{-\frac{\mu^{2}}{\sigma^{2}}+\mathcal{O}\left(\sigma^{2}\right)}-\mu}{\mu}+x\left(\frac{2}{\sqrt{\pi}\sigma}+\mathcal{O}\left(\sigma^{2}\right)\right)e^{-\frac{\mu^{2}}{\sigma^{2}}+\mathcal{O}\left(\sigma^{2}\right)}+\mathcal{O}\left(x^{2}\right)$ (42)
We then calculate the probabilities of no errors ($k=0$) and one logical error ($|k|=1$) giving us the simplified expressions,
$\begin{split}&\tilde{P}_{\mathrm{shift}}(\mathbb{I};\varepsilon)=1-\frac{\sqrt{2d\varepsilon}\exp\left(-\pi/(2d\varepsilon)\right)}{\pi}\\ &\tilde{P}_{\mathrm{shift}}(X^{\pm 1};\varepsilon)=\frac{\sqrt{d\varepsilon}\exp\left(-\pi/(2d\varepsilon)\right)}{\sqrt{2}\pi}.\end{split}$ (43)
@@
$P^{(\mathrm{hex})}(X^{k},\sigma^{2})\equiv P^{(\mathrm{hex})}(Z^{k},\sigma^{2})=\sum_{j\in\mathbb{Z}}\frac{1}{2}\left(\mathrm{erf}\left(\sqrt{\frac{4\pi}{\sqrt{3}d}}\frac{jd+k+1/2}{\sigma}\right)-\mathrm{erf}\left(\sqrt{\frac{4\pi}{\sqrt{3}d}}\frac{jd+k-1/2}{\sigma}\right)\right).$ (44)
@@
$\begin{split}&\tilde{P}^{(\mathrm{hex})}_{\mathrm{shift}}(\mathbb{I};\varepsilon)=1-\frac{\sqrt{\sqrt{3}d\varepsilon}\exp\left(-\pi/(\sqrt{3}d\varepsilon)\right)}{\pi}\\ &\tilde{P}^{(\mathrm{hex})}_{\mathrm{shift}}(X^{\pm 1};\varepsilon)=\frac{\sqrt{\sqrt{3}d\varepsilon}\exp\left(-\pi/(\sqrt{3}d\varepsilon)\right)}{2\pi}.\end{split}$ (45)
@@
We may also analyze the hashing rate as a function of the state’s squeezing. Figs. 6 and 7 show the achievable distillable entanglement rate (absolute and relative scaling respectively) for the dual-homodyne based swap. For a given value of channel loss (panels a-d), we see that the distillable entanglement rate attains a maximum ‘steady state’ value for higher values of squeezing. The upper limit to this rate decreases with increasing channel loss as seen in Fig. 6. Additionally the fraction of the maximum achievable $\mathcal{R}(\tilde{\rho}_{\bm{M}_{1}\bm{M}_{2}})$ (which for a given qudit encoding is $N$) is dependent on the loss and the encoding size. For higher losses, the maximum ‘steady state’ rate fraction (i.e. $\mathcal{R}(\tilde{\rho}_{\bm{M}_{1}\bm{M}_{2}})/N$ ) is lower for larger values of $N$.

### v2

@@ Methods
Assuming ideal GKP qudits, and no loss in transmission, the dual homodyne measurement based Bell state measurement, heralds a pure state |$\ket{\Psi_{k,l}}_{\bm{M}_{1},\bm{M}_{2}}$. If the GKP qudits are transmitted over a link of transmissivity $\eta<1$, thereby undergoing the aforementioned random displacement channel, the final state $\rho_{\bm{M}_{1},\bm{M}_{2}}$ will be a mixed state.
We now consider the approximate models to random displacement. Assuming the pure state evolution noise model, we label the heralded pure state |$\ket{\Psi^{\prime}}_{\bm{M}_{1},\bm{M}_{2}}$. However with the twirled noise model, the final state is mixed and labelled $\tilde{\rho}_{\bm{M}_{1},\bm{M}_{2}}$. The pure state evolution model treats the effect of loss coherently, hence, the entanglement distribution rate estimated by this approach would overestimate the actual performance. One the other hand, the twirled error model ignores residual coherences between the ‘shifted’ states - we ignore terms of the form $\hat{X}^{k_{1}}\hat{Z}^{k_{2}}\ket{\chi}\!\!\bra{\chi}_{G}\hat{Z}^{k^{\prime}_{2}}\hat{X}^{k^{\prime}_{1}}$, where $k_{1}\neq k_{1}^{\prime};k_{2}\neq k_{2}^{\prime}$ in Eq. (39). Hence we expect the following inequality,
$I(\ket{\Psi^{\prime}}\!\!\bra{\Psi^{\prime}}_{\bm{M}_{1},\bm{M}_{2}})\geq I(\rho_{\bm{M}_{1},\bm{M}_{2}})\geq I(\tilde{\rho}_{\bm{M}_{1},\bm{M}_{2}})$ (41)
where $I(\rho_{AB})=\max[S(\rho_{A})-S(\rho_{AB}),S(\rho_{B})-S(\rho_{AB}),0]$ and equality only holds for ideal GKP qudits and lossless transmission ($\eta=1$). Hence, the hashing rate for $\tilde{\rho}_{\bm{M}_{1},\bm{M}_{2}}$ would be an achievable distillable entanglement generation rate for the complete noisy state $\rho_{\bm{M}_{1},\bm{M}_{2}}$ we limit our analysis to $I(\tilde{\rho}_{\bm{M}_{1},\bm{M}_{2}})$ in the main text as well as the Methods.
.4 Hashing Rate Asymptote for Low Loss Transmission with Ideal GKP States
Assuming we start with the state in Eq. (40), where the GKP qudit is an approximate but pure state with variance of $\sigma^{2}$, the probability of making $k$-shift errors in the logical measurement for sq-GKP states (along either the Pauli $X$ or $Z$ basis) is given by [21],
$P^{(\mathrm{sq})}(X^{k},\sigma^{2})\equiv P^{(\mathrm{sq})}(Z^{k},\sigma^{2})$ $=\sum_{j\in\mathbb{Z}}\frac{1}{2}\left(\mathrm{erf}\left(\sqrt{\frac{2\pi}{d}}\frac{jd+k+1/2}{\sigma}\right)-\mathrm{erf}\left(\sqrt{\frac{2\pi}{d}}\frac{jd+k-1/2}{\sigma}\right)\right);$
$\coloneqq{P}^{(\mathrm{sq})}_{\mathrm{shift}}(k;\sigma^{2});\;k\in\{-\lfloor d/2\rfloor+1,-\lfloor d/2\rfloor+2,\ldots,\lfloor d/2\rfloor\},$ (42)
where $k\in\{0,1,2,\ldots,d-1\}$, with $k=0$ denoting the probability of deducing a correct logical outcome, whereas the summation over the index $j$ accounts for shifts of magnitude greater than one lattice spacing. Assuming that Bell state measurement logical outcomes are $l_{Z},l_{X}$ the final state (assuming the twirled noise model) can be expressed as
$\tilde{\rho}_{\bm{M}_{1}\bm{M}_{2}}=\sum_{k_{1},k_{2}=0}^{d-1}P_{\mathrm{shift}}^{(k_{1},k_{2})}(\sigma^{2})\,\ket{\Psi_{k^{\prime},l^{\prime}}}\!\!\bra{\Psi_{k^{\prime},l^{\prime}}}_{\bm{M}_{1}\bm{M}_{2}},$ (43)
where $P_{\mathrm{shift}}^{(k_{1},k_{2};)}(\sigma^{2})=P(X^{k_{1}},\sigma^{2})\cdot P(Z^{k_{2}},\sigma^{2})$ and the state |$\ket{\Psi_{k^{\prime},l^{\prime}}}\equiv W_{1}^{(0,d-l_{Z})}W_{2}^{(d-l_{X},0)}\ket{\Psi_{0,0}}$ is a qudit Bell state. The hashing bound [26, 22] based lower bound to the distillable entanglement for this state is given by
@@
$\tilde{P}^{(\mathrm{sq})}_{\mathrm{shift}}(k;1-\sqrt{\eta})\approx\frac{1}{2}\left(\frac{\left|k+1/2\right|}{k+1/2}-\frac{\left|k-1/2\right|}{k-1/2}+\frac{\sqrt{d\varepsilon}\exp\left(-2\pi(k-1/2)^{2}/(d\varepsilon)\right)}{\sqrt{2}\pi(k-1/2)}-\frac{\sqrt{d\varepsilon}\exp\left(-2\pi(k+1/2)^{2}/(d\varepsilon)\right)}{\sqrt{2}\pi(k+1/2)}\right),$ (45)
@@
$\mathrm{erf}\left(\frac{x-\mu}{\sigma}\right)=\frac{\left(\frac{\sigma}{\sqrt{\pi}}+\mathcal{\mathcal{O}}\left(\sigma^{2}\right)\right)e^{-\frac{\mu^{2}}{\sigma^{2}}+\mathcal{O}\left(\sigma^{2}\right)}-\mu}{\mu}+x\left(\frac{2}{\sqrt{\pi}\sigma}+\mathcal{O}\left(\sigma^{2}\right)\right)e^{-\frac{\mu^{2}}{\sigma^{2}}+\mathcal{O}\left(\sigma^{2}\right)}+\mathcal{O}\left(x^{2}\right),$ (46)
We then calculate the probabilities of no errors ($k=0$) and one logical error ($|k|=1$) giving us the simplified and approximate shift error expressions,
$\begin{split}&\tilde{P}^{(\mathrm{sq})}_{\mathrm{shift}}(0;\varepsilon)=1-\frac{\sqrt{2d\varepsilon}\exp\left(-\pi/(2d\varepsilon)\right)}{\pi};\\ &\tilde{P}^{(\mathrm{sq})}_{\mathrm{shift}}({\pm 1};\varepsilon)=\frac{\sqrt{d\varepsilon}\exp\left(-\pi/(2d\varepsilon)\right)}{\sqrt{2}\pi}.\end{split}$ (47)
@@
$P^{(\mathrm{hex})}(X^{k},\sigma^{2})\equiv P^{(\mathrm{hex})}(Z^{k},\sigma^{2})$ $=\sum_{j\in\mathbb{Z}}\frac{1}{2}\left(\mathrm{erf}\left(\sqrt{\frac{4\pi}{\sqrt{3}d}}\frac{jd+k+1/2}{\sigma}\right)-\mathrm{erf}\left(\sqrt{\frac{4\pi}{\sqrt{3}d}}\frac{jd+k-1/2}{\sigma}\right)\right)$
$\coloneqq{P}^{(\mathrm{hex})}_{\mathrm{shift}}(k;\sigma^{2});\;k\in\{-\lfloor d/2\rfloor+1,-\lfloor d/2\rfloor+2,\ldots,\lfloor d/2\rfloor\},$ (48)
@@
We may also analyze the hashing rate as a function of the state’s squeezing. Figs. 7 and 8 show the achievable distillable entanglement rate (absolute and relative scaling respectively) for the dual-homodyne based swap. For a given value of channel loss (panels a-d), we see that the distillable entanglement rate attains a maximum ‘steady state’ value for higher values of squeezing. The upper limit to this rate decreases with increasing channel loss as seen in Fig. 7. Additionally the fraction of the maximum achievable $\mathcal{R}(\tilde{\rho}_{\bm{M}_{1}\bm{M}_{2}})$ (which for a given qudit encoding is $N$) is dependent on the loss and the encoding size. For higher losses, the maximum ‘steady state’ rate fraction (i.e. $\mathcal{R}(\tilde{\rho}_{\bm{M}_{1}\bm{M}_{2}})/N$ ) is lower for larger values of $N$.